# Safe and Optimal Learning from Preferences via Weighted Temporal Logic with Applications in Robotics and Formula 1

### v1

@@ Abstract
Autonomous systems increasingly rely on human feedback to align their behavior, expressed as pairwise comparisons, rankings, or demonstrations. While existing methods can adapt behaviors, they often fail to guarantee safety in safety-critical domains. We propose a safety-guaranteed, optimal, and efficient approach to solve the learning problem from preferences, rankings, or demonstrations using Weighted Signal Temporal Logic (WSTL). WSTL learning problems, when implemented naively, lead to multi-linear constraints in the weights to be learned. By introducing structural pruning and log-transform procedures, we reduce the problem size and recast the problem as a Mixed-Integer Linear Program while preserving safety guarantees. Experiments on robotic navigation and real-world Formula 1 data demonstrate that the method effectively captures nuanced preferences and models complex task objectives.
@@ I Introduction
Autonomous systems are increasingly part of our daily lives, from driverless cars in urban navigation to household robots performing domestic chores. Since these systems operate closely alongside humans, learning from human feedback is a natural way to ensure their behaviors align with human desires. Humans express their preferences in many forms: pairwise comparisons (“I prefer A over B”) [1], rankings (“I prefer A over B, and B over C”) [2, 3], or demonstrations (examples of trajectories or task executions) [4, 5]. Preference learning and learning-to-rank methods leverage comparative judgments to form the queries and collect preference data about behaviors. These preferences are then used in frameworks such as preference-based learning [6], reinforcement learning from human feedback [7, 8], and direct optimization [9], to tune the system behavior. In learning from demonstrations, users directly demonstrate the task or guide the system, often using tools like teleoperation. After the data collection step, these demonstrations are fed into learning frameworks such as behavioral cloning or inverse reinforcement learning.
@@
While these approaches show success in adapting the system behavior based on preferences, they fall short in providing rigorous safety guarantees when the systems are deployed in safety-critical domains, such as autonomous vehicles, medical robotics, and industrial automation. These methods generally assume that users demonstrate and prefer safe options. However, in cases where the user cannot reliably judge safety, this assumption can be dangerous or even fatal for all agents in the environment. The fundamental challenge is to enable learning within safe behaviors, even in cases where user preferences conflict with safety requirements.
@@
Karagulle et al. offer a solution for the safe preference learning challenge [10, 11]. Combining temporal logic formalism with preference learning using pairwise comparisons, their proposed framework learns the importance order of the sub-tasks and time instances using preference data. More specifically, they express system specifications using Weighted Signal Temporal Logic (WSTL) [12] and learn the weights of the formula. This learned WSTL specification can be used for control synthesis tasks [13]. The framework preserves the qualitative semantics of the task specification, ensuring that the meaning of safety does not depend on learned weights. As WSTL provides a quantitative measure for satisfaction and violation of specifications, the framework guarantees that an unsafe behavior is never favored over a safe one. The authors use gradient-based learning or random sampling for the learning part. While these methods are effective in finding suitable weights, both methods are fundamentally limited in finding the optimal solution in terms of the number of preferences satisfied by the formula. Moreover, gradient-based approaches are susceptible to getting stuck at a local minimum due to the complex nature of the quantitative semantics of the temporal logic. Solving safe preference learning to optimality leads to a Mixed-Integer Problem (MIP) formulation with multi-linear constraints with integers, a hard class of problems to solve.
@@
In this paper, we first extend the safe preference learning problem to richer forms of feedback: pairwise preferences, rankings, and demonstrations. Additionally, we overcome the computational challenges of previous methods by introducing two key procedures. First, structural pruning eliminates parts of the formula that do not impact the quantitative semantics for each signal, thus reducing the problem size. We show that structural pruning does not change the quantitative semantics of signals over formulas. Second, log-transform linearizes temporal logic constraints, thus enabling the formulation of the problem as a Mixed-Integer Linear Program (MILP). We prove that the log-transformation does not affect the optimizer of the optimization problem. Together, these techniques allow us to achieve efficient and optimal learning. This is in contrast to earlier approaches based on heuristic solvers without optimality guarantees [10].
We demonstrate the performance of the proposed method on a safe preference learning problem for a simple robotics navigation task, and a learning-to-rank problem using real-world Formula 1 data. In the robot navigation experiment, we show that the method is sensitive to even small changes in preferences and successfully reflects these differences in the synthesized trajectories. In the Formula 1 experiment, we illustrate that the learned weights capture complex structures like critical aspects of race performance, effectively modeling race strategies and providing insights into what factors influence successful outcomes over the course of the race.
@@
(c) The pruned RCT at $t=0$, with the robustness values noted at edges [^2] We denote $\rho(\sigma,\phi_{i},t^{\prime})$ as $\rho_{\phi_{i}}^{\sigma(t^{\prime})}$ due to space limitations.. Pruned parts are shown with low opacity.
Figure 1: Trees associated with $\phi=\Diamond_{[3,5]}(0\leq\sigma\leq 5)$; AST, RCT, and pruned RCT for signal $\sigma=[5,6,7,-1,4,2]$, respectively.
@@ II-A Signal Temporal Logic (STL)
STL is used to reason about signals $\sigma:\mathbb{T}\to\mathcal{S}$, where $\mathbb{T}\subset\mathbb{Z}_{\geq 0}$ is the time domain and $\mathcal{S}\subseteq\mathbb{R}^{m}$ is the $m$-dimensional real-valued signal domain. A well-formed STL formula $\phi$ follows the grammar $\phi::=\top\mid\mu\mid\lnot\phi\mid\phi_{1}\land\phi_{2}\mid\phi_{1}\text{\bf{U}}_{[a,b]}\phi_{2}$. Syntax elements have the following interpretations: $\top$ is Boolean truth, $\mu$ is a predicate of the form $\mu(\sigma(t)):=h_{\mu}(\sigma(t))\geq 0$ where $h_{\mu}:\mathcal{S}\to\mathbb{R}$ maps the signal value at time instant $t$, to a real value [^3] Note that it is possible to transform any (in)equality of the form $f(x)\sim c$, where $\sim\in\{\leq,\geq,=\}$, into the form defined in the syntax.. The operator $\lnot$ is the negation, $\land$ is the conjunction, and $\text{\bf{U}}_{[a,b]}$ is the “Until” operator [^4] Additional operators; disjunction $\lor$, Always $\square_{[a,b]}$, and Eventually $\Diamond_{[a,b]}$ can be derived from the grammar as $\phi_{1}\lor\phi_{2}=\lnot(\lnot\phi_{1}\land\lnot\phi_{2})$, $\Diamond_{[a,b]}\phi=\top\text{\bf{U}}_{[a,b]}\phi$, and $\square_{[a,b]}\phi=\lnot(\Diamond_{[a,b]}\lnot\phi)$.. Subscript [a,b]{[a,b]} with $a,b\in\mathbb{Z}_{\geq 0}$ defines the time interval that the temporal operator is acting on. When the time interval includes the whole signal length, we omit the time interval subscript. The qualitative semantics of STL are defined in [14]. If a signal $\sigma$ satisfies (violates) a formula $\phi$ at time $t$, it is shown as $(\sigma,t)\models\phi$ ($(\sigma,t)\not\models\phi$). For qualitative semantics at time instant $t=0$, we omit $t$ and write $\sigma\models\phi$. STL has quantitative semantics to measure how well the signal satisfies the formula or how badly it violates the formula. There are different quantitative semantics in the literature [14, 15, 16]. We follow the metric defined in [14], and call it the robustness. The robustness of a signal at time $t$ is shown as $\rho(\sigma,\phi,t)$. When $t=0$, we denote it as $\rho(\sigma,\phi)$. When we work with finite signals, we assume that the signal length covers the time horizon of the formula, a function of time intervals of temporal operators. Additionally, we emphasize that we use STL over discrete-time (as opposed to continuous-time) finite signals with necessary modifications as needed [17].
@@
Each STL formula can be represented by an Abstract Syntax Tree (AST) [18]. In an AST, nodes represent syntactic elements: leaf nodes correspond to predicates or $\top$, while non-leaf nodes correspond to logical and temporal operators. Edges encode the relationships between operators and their operands in each subformula. Unary operators (Negation, Always, Eventually) have one child node, while binary operators (And, Or, Until) have two. To represent and reason about the robustness more systematically, we represent the robustness computation as a tree. Building on the AST, we define the Robustness Computation Tree (RCT) of a formula $\phi$, at time $t$ as $\mathcal{T}_{\phi,t}=(\mathcal{N}_{\phi,t},\mathcal{E}_{\phi,t})$, where $\mathcal{N}_{\phi,t}$ is the set of nodes, obtained by extending each AST node with the explicit time index at which the corresponding operator or predicate is evaluated during robustness computation, and $\mathcal{E}_{\phi,t}$ is the set of directed edges encoding parent–child relationships in the time-indexed structure. Placing the signal’s predicate robustness values $h_{\mu}(\sigma)$ to leaf nodes of $\mathcal{T}_{\phi,t}$, and replacing each operator with its function defined in the quantitative semantics, the tree exactly mirrors the robustness computation for $\rho(\sigma,\phi,t)$, when operations from the leaf nodes to the root node of the tree are followed. Moreover, for any subformula $\phi_{i}$ and $t^{\prime}$ appearing in the formula, the subtree $\mathcal{T}_{\phi_{i},t^{\prime}}$ corresponds to $\rho(\sigma,\phi_{i},t^{\prime})$.
Example 1.
@@
Let $\phi=\Diamond_{[3,5]}(0\leq\sigma\leq 5)$ be an STL formula. Figure 1(a) and 1(b) show the AST and RCT of $\phi$ at $t=0$, respectively. Tree $\mathcal{T}_{0\leq\sigma\leq 5,t^{\prime}}$ corresponds to $\rho(\sigma,0\leq\sigma\leq 5,t^{\prime})$.
@@ II-B Weighted Signal Temporal Logic (WSTL)
WSTL is tailored to represent priorities and preferences in STL formulas [12]. Its syntax extends STL syntax as $\phi::=\top\mid\mu\mid\lnot\phi\mid\phi_{1}\land^{w}\phi_{2}\mid\phi_{1}\text{\bf{U}}_{[a,b]}^{w^{1},w^{2}}\phi_{2}$, where the weights are $w\in\mathbb{R}_{>0}^{2}$ and $w^{1},w^{2}:[a,b]\to\mathbb{R}_{>0}$. All operators are interpreted as in STL. The quantitative semantics of WSTL is called the weighted robustness, and defined as [10]:
[…]
@@
$\begin{array}[]{rcl}r(\sigma,\phi_{1}\lor^{w}\phi_{2},t)&=&\max\big(w_{1}r(\sigma,\phi_{1},t),w_{2}r(\sigma,\phi_{2},t)\big),\\ r(\sigma,\square^{w}\phi,t)&=&\min\limits_{t^{\prime}\in[a,b]}\big(w_{t^{\prime}}r(\sigma,\phi,t+t^{\prime})\big),\\ r(\sigma,\Diamond^{w}\phi,t)&=&\max\limits_{t^{\prime}\in[a,b]}\big(w_{t^{\prime}}r(\sigma,\phi,t+t^{\prime})\big).\end{array}$ (2)
To denote the weighted robustness at $t=0$, we use $r(\sigma,\phi)$. As it is shown in [10, 12], the quantitative semantics in (1) is sound. The RCT can be defined for WSTL formulas as well. Weights of the operators are added as costs to the relevant edges, and we obtain $\mathcal{T}_{\phi_{\boldsymbol{w}},t}=(\mathcal{N}_{\phi,t},\mathcal{E}_{\phi,t},\boldsymbol{w})$.
@@ II-C Parametric Weighted Signal Temporal Logic (PWSTL)
To enable learning from data, we treat weights as parameters, and define an extension to WSTL called Parametric Weighted Signal Temporal Logic (PWSTL) (cf. [19]). We denote the set of unknown parameters as $\mathcal{W}$, and PWSTL formulas as $\phi_{\mathcal{W}}$. The weight set is the set (or a subset) of all possible weights in the formula. A PWSTL formula results in a WSTL formula $\phi_{\mathcal{W}=\boldsymbol{w}}$ with the valuation $\boldsymbol{w}$ of the parameters. For simplicity, if it is clear from the context, we write $\phi_{\boldsymbol{w}}$ to denote the WSTL formula with valuation $\boldsymbol{w}$.

### v2

@@ Abstract
Autonomous systems increasingly rely on human feedback to align their behavior, expressed as pairwise comparisons, rankings, or demonstrations. While existing methods can adapt behaviors, they often fail to guarantee safety in safety-critical domains. We propose a safety-guaranteed, optimal, and efficient approach for solving the learning problem from preferences, rankings, or demonstrations using Weighted Signal Temporal Logic (WSTL). WSTL learning problems, when implemented naively, lead to multi-linear constraints in the weights to be learned. By introducing structural pruning and log-transform procedures, we reduce the problem size and recast it as a Mixed-Integer Linear Program while preserving safety guarantees. Experiments on robotic navigation and real-world Formula 1 data demonstrate that the method captures nuanced preferences and models complex task objectives.
@@ I Introduction
Autonomous systems are increasingly part of our daily lives, from driverless cars in urban navigation to household robots performing domestic chores. Since these systems operate closely alongside humans, learning from human feedback is a natural way to ensure the system’s behaviors align with human desires. Humans express their preferences in many forms: pairwise comparisons (“I prefer A over B”) [1], rankings (“I prefer A over B, and B over C”) [2, 3], or demonstrations (examples of trajectories or task executions) [4, 5]. Preference learning and learning-to-rank methods leverage comparative judgments to form the queries and collect preference data about behaviors. These preferences are then used in frameworks such as preference-based learning [6], reinforcement learning from human feedback [7, 8], and direct optimization [9], to tune the system’s behavior. In learning from demonstrations, users directly demonstrate the task or guide the system, by tools like teleoperation. After data collection, these demonstrations are fed into frameworks such as behavioral cloning or inverse reinforcement learning.
@@
While these approaches have shown promise in adapting the system behavior based on preferences, they fall short in providing rigorous safety guarantees when the systems are deployed in safety-critical domains, such as autonomous vehicles and industrial automation. These methods generally assume that users demonstrate and prefer safe options. However, in cases where the user cannot reliably judge safety, this assumption can be dangerous or even fatal for all agents in the environment. The challenge is to enable learning within the space of safe behaviors, even in cases where user preferences conflict with safety requirements.
@@
Karagulle et al. offer a solution for the safe preference learning challenge [10, 11]. Combining temporal logic formalism with preference learning using pairwise comparisons, their proposed framework learns the importance order of the sub-tasks and time instances using preference data. More specifically, they express system specifications using Weighted Signal Temporal Logic (WSTL) [12] and learn the weights of the formula. This learned WSTL specification can be used for control synthesis tasks [13]. The framework preserves the qualitative semantics of the task specification, ensuring that the meaning of safety does not depend on learned weights. As WSTL provides a quantitative measure for satisfaction and violation of specifications, the framework guarantees that an unsafe behavior is never favored over a safe one. When the weights of the formula are known, WSTL formulas can be encoded as linear constraints with integers [14]. However, when weights are to be learned, encoding yields multi-linear constraints, as weights appear multiplicatively. Thus, solving safe preference learning to optimality leads to a Mixed-Integer Problem with multi-linear constraints, a hard class of problems to solve. Instead, the authors use gradient-based learning or random sampling for the learning part. While these methods are effective in finding suitable weights, both methods are fundamentally limited in finding the optimal solution in terms of the number of preferences satisfied. Moreover, gradient-based approaches are susceptible to getting stuck at local minima due to the quantitative semantics’ complex nature.
@@
In this paper, we first extend the safe preference learning problem to richer forms of feedback: pairwise preferences, rankings, and demonstrations. We overcome the computational challenges with two key procedures. We propose the $\log$-transform to linearize constraints and recast the problem as an MILP, proving it preserves the optimizer. Since the $\log$ function requires strictly positive terms, we introduce structural pruning to eliminate formula parts that do not affect quantitative semantics. Structural pruning, therefore, provides an efficient encoding of constraints as well. We prove that structural pruning preserves the quantitative semantics of signals over formulas. Together, these techniques allow us to achieve efficient and optimal learning, as opposed to heuristic approaches that lack optimality guarantees [10].
@@
(c) Pruned RCT at $t=0$, with the robustness values ($\rho(\sigma,\phi_{i},t^{\prime})$ denoted as $\rho_{\phi_{i}}^{\sigma(t^{\prime})}$) noted at edges. Pruned parts are shown with low opacity.
Figure 1: Trees associated with $\phi=\Diamond_{[3,5]}(0\leq\sigma\leq 5)$; AST, RCT, and pruned RCT for $\sigma=[5,6,7,-1,4,2]$, respectively.
We demonstrate the performance of the proposed method on a safe preference learning problem for a simple robotics navigation task, and a learning-to-rank problem using real-world Formula 1 data. In the robot navigation experiment, we show that the method is responsive to even small changes in preferences and successfully reflects these differences in the synthesized trajectories. In the Formula 1 experiment, we illustrate that the learned weights capture complex structures like race performance, effectively modeling race strategies and providing insights into what factors influence successful outcomes over the course of the race.
@@ II-A Signal Temporal Logic (STL)
STL is used to reason about signals $\sigma:\mathbb{T}\to\mathcal{S}$, where $\mathbb{T}\subset\mathbb{Z}_{\geq 0}$ is the time domain and $\mathcal{S}\subseteq\mathbb{R}^{m}$ is the $m$-dimensional real-valued signal domain. A well-formed STL formula $\phi$ follows the grammar $\phi::=\top\mid\mu\mid\lnot\phi\mid\phi_{1}\land\phi_{2}\mid\phi_{1}\text{\bf{U}}_{[a,b]}\phi_{2}$. Syntax elements have the following interpretations: $\top$ is Boolean truth, $\mu$ is a predicate of the form $\mu(\sigma(t)):=h_{\mu}(\sigma(t))\geq 0$ where $h_{\mu}:\mathcal{S}\to\mathbb{R}$ maps the signal value at time instant $t$, to a real value [^1] Note that it is possible to transform any (in)equality of the form $f(x)\sim c$, where $\sim\in\{\leq,\geq,=\}$, into the form defined in the syntax.. The operator $\lnot$ is the negation, $\land$ is the conjunction, and $\text{\bf{U}}_{[a,b]}$ is the “Until” operator [^2] Additional operators; disjunction $\lor$, Always $\square_{[a,b]}$, and Eventually $\Diamond_{[a,b]}$ can be derived from the grammar as $\phi_{1}\lor\phi_{2}=\lnot(\lnot\phi_{1}\land\lnot\phi_{2})$, $\Diamond_{[a,b]}\phi=\top\text{\bf{U}}_{[a,b]}\phi$, and $\square_{[a,b]}\phi=\lnot(\Diamond_{[a,b]}\lnot\phi)$.. Subscript [a,b]{[a,b]} with $a,b\in\mathbb{Z}_{\geq 0}$ defines the time interval that the temporal operator is acting on. When the time interval includes the whole signal length, we omit the time interval subscript. The qualitative semantics of STL are defined in [15]. If a signal $\sigma$ satisfies (violates) a formula $\phi$ at time $t$, it is shown as $(\sigma,t)\models\phi$ ($(\sigma,t)\not\models\phi$). For qualitative semantics at time $t=0$, we omit $t$ and write $\sigma\models\phi$. STL has quantitative semantics to measure how well the signal satisfies the formula or how badly it violates the formula. There are different quantitative semantics in the literature [15, 16, 17]. We follow the metric defined in [15], and call it the robustness. The robustness of a signal at time $t$ is shown as $\rho(\sigma,\phi,t)$. When $t=0$, we denote it as $\rho(\sigma,\phi)$. When we work with finite signals, we assume that the signal length covers the time horizon of the formula, a function of time intervals of temporal operators. Additionally, we use STL over discrete-time (as opposed to continuous-time) finite signals with necessary modifications as needed [18].
@@
Each STL formula can be represented by an Abstract Syntax Tree (AST) [19]. In an AST, nodes represent syntactic elements: leaf nodes correspond to predicates or $\top$, non-leaf nodes correspond to logical and temporal operators. Edges encode the relationships between operators and their operands in each subformula. Unary operators have one child node, while binary operators have two. To reason about the robustness more systematically, we represent the robustness computation as a tree. Building on the AST, we define the Robustness Computation Tree (RCT) of $\phi$, at time $t$ as $\mathcal{T}_{\phi,t}=(\mathcal{N}_{\phi,t},\mathcal{E}_{\phi,t})$, where $\mathcal{N}_{\phi,t}$ is the set of nodes, obtained by extending each AST node with the time index at which the corresponding operator or predicate is evaluated during robustness computation, and $\mathcal{E}_{\phi,t}$ is the set of directed edges encoding parent–child relationships in the time-indexed structure. Placing the signal’s predicate robustness values $h_{\mu}(\sigma)$ at leaf nodes of $\mathcal{T}_{\phi,t}$, and replacing each operator with its function defined in the quantitative semantics, the tree mirrors the robustness computation for $\rho(\sigma,\phi,t)$ when evaluating from leaf to root. For a subformula $\phi_{i}$ and $t^{\prime}$ of $\phi$, the subtree $\mathcal{T}_{\phi_{i},t^{\prime}}$ corresponds to $\rho(\sigma,\phi_{i},t^{\prime})$.
Example 1
@@
Let $\phi=\Diamond_{[3,5]}(0\leq\sigma\leq 5)$ be an STL formula. Figure 1(a) and 1(b) show the AST and RCT of $\phi$ at $t=0$, respectively. Tree $\mathcal{T}_{0\leq\sigma\leq 5,t^{\prime}}$ represents $\rho(\sigma,0\leq\sigma\leq 5,t^{\prime})$.
@@ II-B Weighted Signal Temporal Logic (WSTL)
WSTL is tailored to represent priorities and preferences in STL formulas [12]. Its syntax extends STL syntax as $\phi::=\top\mid\mu\mid\lnot\phi\mid\phi_{1}\land^{w}\phi_{2}\mid\phi_{1}\text{\bf{U}}_{[a,b]}^{w^{1},w^{2}}\phi_{2}$, where the weights are $w\in\mathbb{R}_{>0}^{2}$ and $w^{1},w^{2}:[a,b]\to\mathbb{R}_{>0}$. All operators are interpreted as in STL. The quantitative semantics of WSTL is called weighted robustness, and defined as [10]:
$r(\sigma,\top,t)$ $=\infty$
$r(\sigma,\mu,t)$ $=\rho(\sigma,\mu,t)$
$r(\sigma,\lnot\phi,t)$ $=-r(\sigma,\phi,t)$
$r(\sigma,\phi_{1}\land^{w}\phi_{2},t)$ $=\min\big(w_{1}r(\sigma,\phi_{1},t),w_{2}r(\sigma,\phi_{2},t)\big)$
$r(\sigma,\phi_{1}\text{\bf{U}}_{[a,b]}^{w^{1},w^{2}}\phi_{2},t)$ $=\hskip-5.69046pt\max\limits_{t^{\prime}\in[a,b]}\hskip-5.69046pt\Big(\min\big(w^{1}_{t^{\prime}}\hskip-5.69046pt\min\limits_{t^{\prime\prime}\in[t,t+t^{\prime})}\hskip-5.69046ptr(\sigma,\phi_{1},t^{\prime\prime}),$
$\quad w^{2}_{t^{\prime}}r(\sigma,\phi_{2},t+t^{\prime})\big)\Big)$ (1)
@@
$r(\sigma,\phi_{1}\lor^{w}\phi_{2},t)$ $=\max\big(w_{1}r(\sigma,\phi_{1},t),w_{2}r(\sigma,\phi_{2},t)\big)$
$r(\sigma,\square^{w}\phi,t)$ $=\min\limits_{t^{\prime}\in[a,b]}\big(w_{t^{\prime}}r(\sigma,\phi,t+t^{\prime})\big)$
$r(\sigma,\Diamond^{w}\phi,t)$ $=\max\limits_{t^{\prime}\in[a,b]}\big(w_{t^{\prime}}r(\sigma,\phi,t+t^{\prime})\big)$ (2)
To denote the weighted robustness at $t=0$, we use $r(\sigma,\phi)$. As shown in [10, 12], the quantitative semantics in (II-B) is sound. The RCT can be defined for WSTL formulas as well. Weights of the operators are added as costs to the relevant edges, and we obtain $\mathcal{T}_{\phi_{\boldsymbol{w}},t}=(\mathcal{N}_{\phi,t},\mathcal{E}_{\phi,t},\boldsymbol{w})$.
@@ II-C Parametric Weighted Signal Temporal Logic (PWSTL)
To enable learning from data, we treat weights as parameters and define an extension to WSTL called Parametric Weighted Signal Temporal Logic (PWSTL) (cf. [20]). We denote the set of unknown parameters as $\mathcal{W}$, and PWSTL formulas as $\phi_{\mathcal{W}}$. The weight set is the set (or a subset) of all possible weights in the formula. A PWSTL formula results in a WSTL formula $\phi_{\mathcal{W}=\boldsymbol{w}}$ with the valuation $\boldsymbol{w}$ of the parameters. For simplicity, if it is clear from the context, we write $\phi_{\boldsymbol{w}}$ to denote the WSTL formula with valuation $\boldsymbol{w}$. Unless noted otherwise, the weight set of the PWSTL version of an STL formula includes all possible weights.